Add. Express answer in simplest form

4/s + 7/s^2
Thank You!!!

Answer by checkley71(8403)

4/S+7/S~2=0 OR (4S+7)/S~2=0 OR 4S+7=0 OR S=-7/4
PROOF 4/(-7/4)+7/(-7/4)~2=0 OR 4*-4/7+7*(-4/7)~2 OR -16/7+7(16/49)=0 OR
-16/7+16/7=0 OR 0=0